### v2

# Did we miss P In CAP? Partial Progress Conjecture under Asynchrony

Abstract
Each application developer desires to provide its users with consistent results and an always-available system despite failures. Boldly, the CALM theorem disagrees. It states that it is hard to design a system that is both consistent and available under network partitions; select at most two out of these three properties. One possible solution is to design coordination-free monotonic applications. However, a majority of real-world applications require coordination. We resolve this dilemma by conjecturing that partial progress is possible under network partitions. This partial progress ensures the system appears responsive to a subset of clients and achieves non-zero throughput during failures. To this extent, we present the design of our Cassandra consensus protocol that allows partitioned replicas to order client requests.
I Introduction
More than two decades ago, when Brewer proposed the CAP theorem, it became evident to database practitioners and researchers that it is hard to design a system that is both consistent and available under network partitions [1, 2, 3]. Unsurprisingly, each application designer still wants to offer its client consistent results and an always-available system despite any failures [4, 5].
Node (or server) failure is the most common type of failure. To handle node failures (availability), application developers employ replication by deploying multiple replicas. Moreover, an application’s clients are often spread across the globe, which results in these replicas being distributed across the globe for low latency to the clients. Having multiple replicas necessitates keeping these replicas in sync (consistency). To do so, these applications run consensus protocols like Paxos [6] and PBFT [7], which aim to establish an agreement among the replicas on the order of executing client transactions despite replica failures.
Unfortunately, replica failures are not the only common type of failure; the more catastrophic ones are network partitions, which hamper communication among the replicas, and thus consensus [8]. As a result, applications suffering from network partitioning are unable to make progress [9, 10]. One way to ensure that an application is both consistent and available under network partitions is if it adheres to the CALM theorem, which expects applications to be monotonic and lack any coordination [11, 12]. However, a majority of real-world applications require coordination and are often non-monotonic (refer to Figure 1).
This makes us wonder, if there exists a mechanism to allow non-CALM applications make progress under network partitions. In this paper, we show that under partitions, partial progress is possible for non-CALM applications. Specifically, we conjecture that the “P” in CAP theorem could be broadened to capture partial progress instead of partition. This partial progress guarantees that the application appears responsive to a subset of clients and achieves a non-zero throughput. Moreover, recovering from these failures requires replicas to run consensus on the pending requests, which can bottleneck the replicas and can lead to catastrophic failures [13].
To allow an application to make partial progress under network partitions, we present the design of a new fault-tolerant paradigm, Cassandra. Like existing CFT protocols [6, 14] and BFT [7], Cassandra also requires a proposal selection phase and proposal commitment phase. However, Cassandra requires each replica to determine the strongest proposal among all the proposals and vote to support this strongest proposal.
Figure 1: CAP vs. CALM applications.
Determining the strongest proposal is a two-step process: (1) comparing rank of the proposals, and (2) comparing the suffix of the logs of committed transactions. Each proposer runs a ranking function to assign its proposal a rank. Each replica marks a proposal as the strongest if its proposer has the most committed transaction. If two proposer’s have equal set of committed transactions, select the proposal with the highest rank. This two-step process of Cassandra allows replicas to continue processing client requests during network partitions and merge their states once the network is synchronous. On merging the states, if any conflict arises, Cassandra provides simple rules to reconcile these conflicts in such a manner that at least one subset of client requests are preserved.
This paper aims to not only present Cassandra but to initiate a discussion that partial progress is possible under network partitions. We show that Cassandra can be adopted by existing applications to achieve non-zero throughput under failures. Although Cassandra follows the design of traditional Paxos protocol, the ideas presented in this paper can be easily extended to modern CFT protocols [14, 15, 16]. Similarly, one can attempt to extend Cassandra to Byzantine failures, as recovering from Byzantine failures is more expensive as shown by literature on Byzantine-fault tolerant systems [17, 18, 19, 20, 21, 22, 23].
To illustrate our partial progress conjecture in practice, we deploy our Cassandra protocol in the wild over the Twitter data. We aim to capture the impacts of delays introduced by network partitions on social interactions such as tweeting (with mentions) and re-tweeting along with quoting, replying, and liking a tweet. Our results show that once the network is restored, the system throughput peeks at $6\times$ of the steady state due to partial progress made during the partition.
II Cassandra at a Glance
From a bird’s-eye view, Cassandra looks like Paxos. Like Paxos, Cassandra provides following desirable properties:
  • Equal Leadership Opportunity. There is no pre-designated leader; each replica can send a proposal.
  • Two-phase Linear Consensus. Each replica runs a two-phase protocol to accept a proposal.
Additionally, Cassandra provides following new properties:
  • Partial Progress under Network Partition. If replicas of a system are unable to communicate with each due to a partitioned or asynchronous network, Cassandra allows these replicas to speculatively order and execute client transactions; Cassandra guarantees that transactions from at least one partition will persist.
  • Multi-Proposal Acceptance. Cassandra supports partial-ordering, which allows it to accept non-conflicting proposals from multiple replicas.
II-A Preliminaries
Prior to explaining the design of Cassandra, we lay down the system model. We make standard assumptions also made by existing CFT protocols [6, 14].
We assume a replicated system $\mathcal{S}=\{\mathcal{R},\mathcal{C}\}$, where $\mathcal{R}$ denotes the set of replicas and $\mathcal{C}$ denotes the set of clients. This replicated system has a total of $\mathbf{n}$ replicas, of which at most $\mathbf{f}$ are faulty; $\mathbf{n}=2\mathbf{f}+1$. The $\mathbf{f}{}$ faulty replicas can fail stop or crash; we do not assume any Byzantine failures. We adopt the same partial synchrony model as existing CFT protocols; safety is guaranteed under asynchrony, while liveness is only guaranteed during periods of synchrony.
$R_{3}$$R_{2}$$R_{1}$c{c}{}T{T}ProposeVotePrepareVotePrepCommit
Figure 2: Schematic representation of the failure-free flow of Cassandra. We assume that $R_{2}$ has the strongest proposal.
II-B Failure-Free Flow
A system is network partitioned if its replicas are divided into groups or partitions such that these partitions cannot exchange messages. In such a setting, the system makes “partial progress” if the requests ordered by at least one partition commit once the network is synchronous. Partial progress under asynchrony / network partitions illustrates that the system processes a subset of client requests and yields a non-zero throughput. Cassandra is the first consensus protocol to guarantee partial progress.
Cassandra ensures partial progress by determining the strongest proposer for each round of consensus. Finding the strongest proposer is a two-step process: (1) comparing rank of the proposals, and (2) comparing the suffix of the logs of committed transactions. Each proposer locally runs a ranking function to assign its proposal a rank; higher the rank, greater are the chances of a proposal being selected. The proposer sends the rank of its proposal along with the proposal. Additionally, the proposer piggybacks its log of transactions. A replica’s log of transactions states the transactions it has voted, prepared, and committed, in order. Cassandra always selects the proposer with a larger log of committed transactions as the strongest proposer.
In Figure 2, we schematically represent our Cassandra protocol. Cassandra requires two phases to achieve consensus among its replicas. The first phase aims to select the strongest proposer (§II-C), while the second phase aims to commit the selected proposal. Cassandra makes use of timers in the first phase for rapid convergence (§II-D); Cassandra is safe and live without timers. Each replica maintains an ordered log of transactions (denoted as histhist). Next, we describe Cassandra in detail and present its pseudocode in Figure 3.
Proposal. The first phase of Cassandra starts when the $i$-th replica $R_{i}$ has a client transaction T{T} that it wants to propose (Line 1). Next, $R_{i}$ creates a proposal $\phi_{i}$, which it broadcasts to all the replicas (Lines 2-5). This proposal includes (1) hash of the ordered log histhist, (2) client transaction and its hash, and (3) a function $\mathop{\textsf{DRF}}()$ that assigns a rank to $R_{i}$’s proposal (§II-C). Post this, $R_{i}$ sets its proposal as the strongest proposal ($Max$) and waits on a timer ($\tau_{p}$) to receive all the other proposals.
Proposal Selection. While waiting on the timer $\tau_{p}$, replica $R_{i}$ may receive proposals (Line 8) from several replicas. It uses these proposals to determine the strongest proposer. It compares the ordered log (histhist) of each incoming proposal against $Max$ and sets $Max$ to the stronger proposal if any. As described above, if two proposals have same ordered logs, then $R_{i}$ updates $Max$ to the proposal with higher rank (Line 12-14). Instead, if a received proposal’s log has larger suffix, we update $Max$ to that proposal.
This implies that a replica with a lower rank but larger suffix gets higher preference than the replica with higher rank and smaller suffix. We explain this choice of update in Section II-E.
Prepare. When replica $R_{i}$’s timer $\tau_{p}$ timeouts, it sends a message Vote to the replica which has been selected as strongest proposer $Max$. Following this, each replica waits to receive the votes. Once a replica $R_{i}$ receives $\mathbf{f}{}+1$ matching Vote messages, it assumes that its proposal has been accepted by a majority of replicas. Post this, $R_{i}$ asks all the other replicas to prepare themselves by broadcasting a Prepare message.
Commit. When a replica $R_{i}$ receives a Prepare message from the $j$-th replica $R_{j}$, it acknowledges this message by sending VotePrep message. Once the replica $R_{j}$ receives identical VotePrep message from $\mathbf{f}{}+1$ replicas, it adds T{T} to its log histhist and broadcasts Commit message. This Commit message allows all the replicas to eventually commit the transaction.
0:Initialization:
// $\mathop{\textsf{DRF}}(i)$ determines strength of $i$-th replica’s proposal.// histhist is the ordered log of transactions.// $Max$ is the strongest proposal for a round.
0:
0:Replica-role (used by $i$-th replica $R_{i}$) :
1:event Received a client transaction T{T} to propose do
2:  Compute hash of ordered log $\mathcal{L}_{i}:=\texttt{Hash}(hist)$.
3:  Compute hash of transaction $\Delta_{i}:=\texttt{Hash}({T})$.
4:  Create proposal $\phi_{i}:=\textsc{Propose}(\mathop{\textsf{DRF}}(i),\mathcal{L}_{i},{T},\Delta_{i})$.
5:  Broadcast $\phi_{i}$ to all the replicas.
6:  Set $Max:=\mathop{\textsf{DRF}}(i)$.
7:  Start a timer $\tau_{p}$
7:
8:event Received a proposal $\phi_{j}$ from $j$-th replica $R_{j}$ do
9:  Compare the suffixes of the two logs.
10:  if $\mathcal{L}_{j}\supset\mathcal{L}_{i}$ then
11:   $Max:=\mathop{\textsf{DRF}}(i)$
12:  else if $\mathcal{L}_{j}=\mathcal{L}_{i}$ then
13:   if $\mathop{\textsf{DRF}}(j)>Max$ then
14:    $Max:=\mathop{\textsf{DRF}}(i)$
14:
15:event Timer $\tau_{p}$ timeouts do
16:  Send $\textsc{Vote}(i,\Delta_{Max})$ to replica with MaxMax proposal.
17:event Received $\mathbf{f}{}+1$ matching $\textsc{Vote}(i,\Delta_{Max})$ messages do
18:  Broadcast $\textsc{Prepare}(i,\Delta_{Max})$
18:
19:event Received $\textsc{Precommit}(i,\Delta_{Max})$ from $j$-th replica do
20:  Broadcast $\textsc{VotePrep}(i,\Delta_{j})$
21:event Received $\mathbf{f}{}+1$ matching $\textsc{VotePrep}(i,\Delta_{j})$ messages do
22:  Add T{T} to ordered log histhist.
23:  Broadcast $\textsc{Commit}(i,\Delta_{j})$
24:event Received $\textsc{Commit}(i,\Delta_{j})$ from $j$-th replica do
25:  Add T{T} to ordered log histhist.
Figure 3: Cassandra protocol (failure-free path).
II-C Deterministic Ranking Function
A key to success of Cassandra lies in determining the strongest proposer among all the proposers. Identifying the strongest proposal helps to converge the consensus faster as all the replicas would vote in the favor of such a proposal. Doing so, however, faces the following challenge: the strength function should be fair and should guarantee that only one proposal is marked as strongest in each round of consensus.
Existing literature presents several ways, which can be used to design such a strength function: (1) Using a verifiable random function (VRF) to generate random numbers that have extremely low probability of collisions [24]. (2) Proposal with highest priority or fees [25], (3) Proposal that solves a puzzle fastest [26, 27], and so on. Although all of these designs yield desirable strength functions, they suffer from real-world adoption as no organization would like to pay for large compute and time. As a result, in this paper, we design a simple, yet efficient deterministic ranking function (DRF).
We use DRF to assign each replica a unique rank in the range $\bf{[0,\mathbf{n}{}]}$. Each replica calls the function $\mathop{\textsf{DRF}}()$ to determine the rank of a replica. The system administrator can decide the frequency of updating ranks of replicas, at the end of each consensus or post some interval of time. In Cassandra, higher the rank of a replica, higher is the chance of its proposal being accepted.
Example 1.
Assume a system of $\mathbf{n}{}=3$ replicas, where ranks of replicas $R_{1}$, $R_{2}$, and $R_{3}$ is $0,2,$ and $1$, respectively. Each replica broadcasts a new proposal during the proposal phase. During the proposal selection phase, each replica receives proposals from all the replicas. As $R_{2}$ has the highest rank, each replica will select the proposal of replica $R_{2}$ as the $Max$ proposal, which will ensure that $R_{2}$ receives the necessary $\mathbf{f}{}+1$ votes.
II-D Rapid Convergence
A CFT consensus protocol is only beneficial if it can guarantee eventual agreement of replicas on a single proposal. In Lamport’s Paxos [6] protocol, it is hard to guarantee that replicas will ever converge, and thus, it cannot guarantee liveness. Recent works [14, 15] eliminate this challenge by designating one replica as the leader. As soon as a replica receives a proposal from the leader, it prepares itself for committing the leader’s proposal. Although designating one replica as the leader is the easiest solution, it suffers three challenges: (1) It lacks fairness. (2) It requires detecting and replacing the leader once it fails, which is not only expensive but hurts system throughput as no new transactions can be ordered until the new leader ensures that all the replicas have the same state. (3) It prevents the system from making partial progress during a network partition if the leader is partitioned from the clients.
Due to these reasons, Cassandra avoids designating any replica as the leader. Instead, it employs DRFs to assign each replica a rank, which helps in determining the strength of a proposal. Cassandra allows the rank of a replica to change over time; we can require each replica to run the DRF before each round of consensus. Although DRFs guarantee fairness, they cannot eliminate the other two challenges. We still need to provide uninterrupted transaction ordering under failures and network partitions. We argue that the use of timers is necessary for uninterrupted transaction processing. We illustrate this through the following schemes, which do not require replicas to wait on a timer (assume that Line 7 in Figure 3 did not exist). Allow of these schemes try to converge the replicas to a single proposal.
  • Attempt 1. We can ask all replicas to wait till they hear the proposal from the replica with the highest rank. Such a condition can cause a replica to hold off voting until it has seen all the proposals as the replica with strongest proposal may be slowest to broadcast or is suffering from message delays. Worse, messages from the replica with the strongest proposal never arrive as it has crashed or partitioned, which will cause the system to get stuck.
  • Attempt 2. We can ask each replica to vote for the first proposal that it receives. Such a solution has extremely low probability of success because replicas are often spread across the globe to guard against data-center failures. Unless at least $\mathbf{f}{}+1$ replicas receive the proposal from the replica with the strongest proposal prior to any other proposal, this solution will not lead to convergence.
  • Attempt 3. We can ask a replica to vote for a proposal, once it has received proposals from a majority of replicas (in our case, $\mathbf{f}{}+1$). Although this solution is better than Attempt 2, it still has a high probability of non-convergence as the first $\mathbf{f}{}+1$ proposals for a majority of replicas may not include the proposal from the replica with strongest proposal.
The non-convergence of these attempts makes us settle down for timers. Each replica initiates a timer after it receives a proposal, and once its timer expires, it compares the DRFs of these proposals to select the strongest proposal. Although not all proposals may arrive until timeout, the system has a flexibility to tune the timeout value. If replicas observe that the timeout period is large (all the proposals are arriving way earlier) and is increasing the latency of convergence, it can propose decreasing the timeout value. Instead, replicas can propose increasing the timeout value if they are unable to converge on a proposal.
II-E Determining Strongest Proposal
One of the goals of Cassandra, like most CFT protocols, is to ensure that all the replicas agree to a common order for all the client transactions. This common ordering choice impacts the way we can select a replica as the strongest proposer. Just simply because a replicas has strongest DRF does not guarantee that it has the strongest proposal. Consider the following example.
Example 2.
Assume a system of $\mathbf{n}{}=3$ replicas $R_{1}$, $R_{2}$, and $R_{3}$, such that all the replicas have committed and logged transaction ${T}_{0}$. Additionally, $R_{2}$ has committed ${T}_{1}$ and prepared ${T}_{2}$, while $R_{3}$ has committed both ${T}_{1}$ and ${T}_{2}$. Now, all the replicas are proposing for the subsequent round and assume the DRF sets the ranks for $R_{1}$, $R_{2}$, and $R_{3}$, as $3,2,$ and $1$, respectively. Despite having the strongest DRF, the proposal of $R_{1}$ never gets $\mathbf{f}{}+1$ votes as its log is missing entries for ${T}_{2}$ and ${T}_{3}$. Although $R_{3}$ has committed ${T}_{2}$, $R_{2}$’s proposal gets $\mathbf{f}{}+1$ votes as it has both prepared ${T}_{2}$ and has a stronger DRF.
This example illustrates that selecting the strongest proposer requires analyzing the rank and suffix of the log of each proposer. Given two replicas $R_{i}$ and $R_{j}$ with logs $hist_{i}$ and $hist_{j}$ and $\mathop{\textsf{DRF}}(i)$ and $\mathop{\textsf{DRF}}(j)$, respectively, we generalize replica selection scheme (for setting $Max$ in Figure 3) as follows:
  1. 1.
    If both the logs have same suffix, then $Max$ is set to the replica with the strongest DRF.
  2. 2.
    If suffix of $hist_{i}$ is $\{{T}_{x}\}$ and $hist_{j}$ is $\{{T}_{x},Prep({T}_{y})\}$, such that $x<y$, then $Max=\mathop{\textsf{DRF}}(j)$.
  3. 3.
    If suffix of $hist_{i}$ is $\{{T}_{x}\}$ and $hist_{j}$ is $\{{T}_{x},{T}_{y}\}$, such that $x<y$, then $Max=\mathop{\textsf{DRF}}(j)$
  4. 4.
    If suffix of $hist_{i}$ is $\{{T}_{x},Prep({T}_{y})\}$ and $hist_{j}$ is $\{{T}_{x},{T}_{y}\}$, such that $x<y$, then $Max$ is set to the replica with the strongest DRF.
In these rules, $Prep({T}_{y})$ implies that a replica has prepared transaction ${T}_{y}$, but has not received Commit message for ${T}_{y}$.
A keen reader would have observed that we require replicas to compare the suffixes of two logs when determining the stronger proposal. Just comparing the suffixes is sufficient due to the following two reasons: (1) Initially, each replica has the same state, and (2) The strongest proposer only marks its proposal as committed if it receives VotePrep messages from a majority ($\mathbf{f}{}+1$) of replicas. Once the strongest proposer marks its proposal as committed, there is a guarantee that in every set of $\mathbf{f}{}+1$ replicas in the system, there will be one replica that has committed the proposal.
III Network Partitions and Failures
It is common for replicas to crash and get network partitioned [9]. These failures not only impact a system’s responsiveness to its clients but also have the potential of making the system stuck. Before we illustrate how Cassandra deals with failures, we try to classify all the failures that prevent consensus among the replicas.
  • Strongest Proposer Failure. The replica with the strongest proposal may fail at any time after it has been determined as the strongest proposer (receives $\mathbf{f}{}+1$ votes). This could lead to replicas being in distinct states with respect to the strongest proposal, and Cassandra needs to bring all these replicas to the common state.
    For instance, consider the following worse-case scenario where the replica $R_{i}$ with the strongest proposal $\phi_{i}$ fails unexpectedly. The timing of this failure can divide replicas into four non-empty sets. Set $A$ includes less than $\mathbf{f}{}$ replicas that did not receive proposal $\phi_{i}$. Set $B$ includes more than $\mathbf{f}{}+1$ replicas that voted for $\phi_{i}$. Set $C$ includes at least $\mathbf{f}{}+1$ replicas that have prepared $\phi_{i}$. Set $D$ includes at most $\mathbf{f}{}$ replicas that have committed $\phi_{i}$.
  • Replica Failure. It is possible that a non-strongest proposer replica fails any time during consensus. This failing replica may be the only replica (apart from the strongest proposer) that has committed the strongest proposal. In such a case, Cassandra needs to ensure that the committed proposal persists despite failures.
  • Replica Partitioning. If the system starts experiencing network partitions, then its replicas may not be able to communicate with each other. A network partitioning could lead to replicas being partitioned (divided) into two or more groups. Following this, the system can be in a state where either no group has at least $\mathbf{f}{}+1$ replicas, or exactly one group has $\mathbf{f}{}+1$ replicas.
Failure Detection. Cassandra aims to quickly detect failures, so that it can run recovery procedures. To detect failures, in each round of consensus, we require each replica to maintain two additional timers: vote timer and prepare timer. Each of these timers start at the end of the preceding phase and are used as follows: (1) Each replica $R_{i}$ starts a vote timer once it sends a Vote message in support of the strongest proposal for that round. When $R_{i}$ receives a Prepare message, it resets the vote timer. (2) Each replica $R_{i}$ starts a prepare timer once it sends a VotePrep message. $R_{i}$ resets this prepare timer after it receives a Commit message.
If either of these timers expire, $R_{i}$ detects a failure and runs the Merge procedure, which we explain next.
III-A Merge Procedure
Cassandra’s novel merge procedure aims to help replicas recover from the various types of failures described earlier. It performs the following tasks: (1) prevent committed state from being lost, and (2) bring all the replicas to the common state. As it is impossible to detect whether a replica has failed or it has partitioned (in an asynchronous network) our merge protocol provides a single algorithm to recover from these failures. This requires us to slightly modify the rules that determine the strongest proposal (§II-E).
Modified Determining Strongest Proposal. These modified rules for determining a strongest proposal allow additional entries in the suffix of a log. Specifically, a replica’s log can include transactions that have neither prepared or committed. Later in this section, we explain why we allow adding such entries to the log. For now, we denote these unprepared and uncommitted transactions as ellipsis ($\dots$) in the log. Given two replicas $R_{i}$ and $R_{j}$ with logs $hist_{i}$ and $hist_{j}$ and $\mathop{\textsf{DRF}}(i)$ and $\mathop{\textsf{DRF}}(j)$, respectively, the strongest proposer ($Max$ in Figure 3) is set as follows:
  1. 1.
    If suffix of $hist_{i}$ is $\{{T}_{x},$...\} and $hist_{j}$ is $\{{T}_{x},$...\}, then $Max$ is set to the replica with the strongest DRF.
  2. 2.
    If suffix of $hist_{i}$ is $\{{T}_{x},$...\} and $hist_{j}$ is $\{{T}_{x},...,Prep({T}_{y}),$...\}, such that $x<y$, then $Max=\mathop{\textsf{DRF}}(j)$.
  3. 3.
    If suffix of $hist_{i}$ is $\{{T}_{x},$...\} and $hist_{j}$ is $\{{T}_{x},....,{T}_{y},$...\}, such that $x<y$, then $Max=\mathop{\textsf{DRF}}(j)$
  4. 4.
    If suffix of $hist_{i}$ is $\{{T}_{x},....,Prep({T}_{y}),$...\} and suffix of $hist_{j}$ is $\{{T}_{x},...,{T}_{y},$...\}, such that $x<y$, then $Max$ is set to the replica with the strongest DRF.
  5. 5.
[…]
  6. 6.
    If suffix of $hist_{i}$ is $\{{T}_{x},....,Prep({T}_{y}),$...\} and suffix of $hist_{j}$ is $\{{T}_{x},...,{T}_{z},$...\}, such that $x<y\wedge x<z$, then $Max=\mathop{\textsf{DRF}}(j)$.
For instance, the first rule states the following: if the last transaction committed by two replicas is the same, then despite any set of unprepared transactions in the log of these replicas, the replica with the strongest DRF is selected as the strongest proposer. A similar interpretation applies to the other rules.
The first four rules are straightforward extension of rules defined in Section II-E. The last two rules are new and help in conflict resolution. The need for conflict resolution occurs when at least one replica has only prepared a transaction ${T}_{y}$, while another replica has instead prepared/committed a transaction ${T}_{z}$. For instance, such a situation can arise if the network is getting repartitioned. In this situation, a replica $R_{i}$ only receives Prepare messages for ${T}_{x}$ (no Commit message) while another replica $R_{i}$ never prepares ${T}_{x}$ (as it did not receive Prepare messages for ${T}_{x}$) but commits ${T}_{y}$ as it receives both Prepare and Commit messages for ${T}_{y}$.
III-A1 Protocol Steps
Next, we explain the Merge procedure that uses the aforementioned strongest proposal rules to allow system to make partial progress under failures.
Vote Timeout. If a replica $R_{i}$’s vote timer timeouts while waiting for the Prepare message, it assumes that the replica it voted as the strongest proposer has failed. As a result, it abandons this round (say $r$) of consensus and ignores any proposals from other replicas for round $r$. As $R_{i}$ was never selected as the strongest proposer in round $r$, it adds its own proposal for round $r$ to its log histhist and starts a new round of consensus ($r+1$). Notice that $R_{i}$ marks its proposal as unprepared in the log. In round $r+1$, once $R_{i}$ has a new transaction to propose, it constructs a new proposal that extends its log histhist and follows the remaining steps in Figure 3.
Prepare Timeout. If a replica $R_{i}$’s prepare timer timeouts while waiting for the Commit message, it assumes that the strongest proposer has failed after sending a Prepare message. As a result, $R_{i}$ adds this strongest proposal to its log and marks its as prepared in its log. Next, $R_{i}$ terminates this round (say $r$) of consensus. When $R_{i}$ has a new transaction to propose, it constructs a new proposal that extends its log histhist and follows the remaining steps in Figure 3 to start a new round of consensus.
Conflict Resolution. During the proposal selection phase (§II-B) a replica $R_{i}$ may receive a conflicting proposal $\phi_{j}$ from a replica $R_{j}$. This conflict is the result of $R_{i}$$R_{j}$ preparing distinct transactions. Fortunately, our rules for determining the strongest proposal dictate which replica’s proposal should be considered strongest. If $R_{i}$’s proposal is weaker, it rollbacks its state.
State Exchange. During the proposal selection phase (§II-B) a replica $R_{i}$ may receive a proposal $\phi_{j}$ from a replica $R_{j}$ such that the suffix of log of $\phi_{j}$ ($hist_{j}$) is greater than the suffix of the log of current strongest proposal $Max$. As per our aforementioned rules, $R_{i}$ should set $Max$ to $\phi_{j}$. This implies that $R_{i}$ is ready to accept $R_{j}$’s log as the most up-to-date history. As $R_{i}$’s log is missing some of the transactions, it queries $R_{j}$ for these missing entries.
IV Evaluation
To evaluate our partial progress conjecture in practice, we deploy our Cassandra protocol in the wild over the Twitter data. We aim to capture social interactions such as tweeting (with mentions) and re-tweeting along with quoting, replying, and liking a tweet, all modeled with transactional semantics. More importantly, we further evaluate the impacts of delays introduced by network partitions that may occur due to possible adversarial attacks in order to disrupt the flow of information or by injecting misinformation (i.e., conflicting transactions) into a partitioned social network. Our aim is to study the effects on throughout and latency upon network recovery to demonstrate partial progress as the system heals itself.
Figure 4: Impacts of network partition on throughput and latency when partial progress is enabled.
Setup and Benchmark. We run our experiments on a local cluster and deploy each replica on a virtual machine having a $4$-core CPU and $8\text{\,}\mathrm{GB}\text{/}$ memory. We run Cassandra on $100$ replicated machines, where the primary mode of communication is gossip to exhibit decentralized deployment. We inject 10,000 transactional tweets per second, where each transaction has a payload of under 1KB. The experiments are run for 260 seconds, where at the 60 seconds mark, the network undergoes partition, the partition lasts for 60 seconds, and the network is restored at 120 seconds. To collect results after reaching a steady state, we discard the measurement during the warmup period, and measurement results are collected over three runs.
Promise of Partial Progress. We evaluate the throughput and latency of the Cassandra in partitioned setting where no partition has the majority of replicas. We observe during the partition, Cassandra enters a speculative ordering and execution phase, where transactions are only softly committed and finality is only established once the partitions are recovered. Therefore, once the network is restored, we observe throughout peeks at $6\times$ of the steady state made possible by enabling partial progress. As expected, the finality latency is delayed proportionally to the length of the network outage. The results are shown in Figure 4.
V Related
Brewer’s CAP theorem [1] argues that while designing an application, the application designer can only select two out of the following three properties: consistency, availability and partition tolerance. Following this, several new models have appeared, which present fresh perspectives for an application designer. PACELC [28] captures the following double trade-off: if partitions occur, then trade between availability and consistency; else trade between latency and consistency. FIT [29] re-imagines the problem of a partitioned system as the problem of fairness and envisions fairness as a metric for latency. CAC [30] redefines consistency as causal consistency and compares it against availability and convergence. BASE [31] proposes a diametrically opposite model to ACID semantics. Unlike pessimistic nature of ACID transactions, BASE vouches for optimistic execution and accepts that the database consistency will be in a state of flux. CALM [11] proponates the ideas proposed by Conflict-replicated data types (CRDTs) [32]. Prior works on designing fault-tolerant database systems include Google’s Spanner [33] and Sequoia [34], which have been adopted in practice. None of these works discuss the possibility of partial progress under CAP model, which is the theme of this paper, and we meet this aim with the help of our Cassandra protocol.
VI Conclusion
In this paper, we illustrated that it is possible to design a system that is consistent, available, and makes partial progress under network partitions. We claim that this partial progress can help an application be responsive to a subset of its clients and allow it to achieve non-zero throughput. To prove our claim, we present the design of our novel Cassandra protocol that allows replicas to continue ordering client transactions even under failures. Cassandra achieves this goal by requiring each replica to vote for the strongest proposal among all the proposals. Selecting the strongest proposal requires comparing the rank of a proposal and observing the suffix of its log. Our experiments demonstrate that Cassandra works in the wild and the partial progress during network partition helps to yield high throughputs once the network is restored.